OpenStudy (mkmkasim):

∫ (x+4)/(x^2 +2x+5) dx

OpenStudy (fibonaccichick666):

Can you factor the bottom at all?

OpenStudy (anonymous):

no, it has no real zeros

OpenStudy (mkmkasim):

so would you have to divide x+4 by the denominator?

OpenStudy (anonymous):

i would complete the square on the bottom, then use a trig sub, that might work

OpenStudy (anonymous):

you cant use partial fractions because the denominator has not real roots, unless you want to be fancy and use partial fractions with complex numbers i would not go that route

OpenStudy (mkmkasim):

yeah thats where i got stuck...

OpenStudy (anonymous):

denominator can be written as \((x+1)^2+4\) which looks very much like \(u^2+a^2\) with \(u=x+1,b=2\)

OpenStudy (fibonaccichick666):

ok so, let's make that bottom into something we can use

OpenStudy (anonymous):

set up for arctangent

OpenStudy (mkmkasim):

ok can you go through how you got to arctangent...

OpenStudy (fibonaccichick666):

So, take a look at the bottom. We want to factor that into something we can use. Start by completing the square

OpenStudy (anonymous):

\[\int\frac{du}{u^2+a^2}=\frac{1}{a}\tan^{-1}(\frac{u}{a})\]

OpenStudy (anonymous):

break in to two parts use a u-sub for the first part, and arctangent for the second

OpenStudy (anonymous):

@FibonacciChick666 you can message me now i think it is ok once you use a gimmick, as all of these require a gimmick of some sort

OpenStudy (anonymous):

if it was \[\int\frac{x+1}{x^2+2x+5}dx\]you could use a simple u - sub \(u=x^2+2x+5, du =(2x+2)dx,\frac{1}{2}du=(x+1)dx\) etc then it would be easy

OpenStudy (anonymous):

since it is not \[\int\frac{x+1}{x^2+2x+5}dx\] but rather \[\int\frac{x+4}{x^2+2x+5}dx\] gimmick is to break in to two parts \[\int\frac{x+1}{x^2+2x+5}dx+\int\frac{3}{x^2+2x+5}dx\]

OpenStudy (anonymous):

first is the simple u - sub as above, you can pretty much do it in your head second is the complete the square get the arctangent part

OpenStudy (mkmkasim):

ok so after you manipulate it with the trig sub, would you eventually get to something like 1/2∫x+4 ?

OpenStudy (fibonaccichick666):

not quite.

OpenStudy (mkmkasim):

ok where did I go wrong, i had the denominator being 4(sec^x), and that canceled out with the du = 2 (secx)^2.

OpenStudy (fibonaccichick666):

It's not a substitution in the form that you are thinking. it is just making it look like something we know how to integrate

OpenStudy (mkmkasim):

ok so how would that look like...

OpenStudy (fibonaccichick666):

Alright let's walk through step by step. Now, as noted by satelite above, the bottom turns into a pretty \((x+1)^2+4\)

OpenStudy (fibonaccichick666):

We want to make the top of our fraction something we can use.

OpenStudy (fibonaccichick666):

So notice that when we make our denominator pretty, it almost looks like a special case we know

OpenStudy (fibonaccichick666):

Do you see what I'm referring to?

OpenStudy (mkmkasim):

ummm maybe.... i know that u^2 +a^2 allows you to use trig sub, but apparently that was wrong...

OpenStudy (fibonaccichick666):

ok, so what is the derivative of \(\frac{1}{a}tan^{-1}(\frac{u}{a})\)

OpenStudy (mkmkasim):

1/1+a^2 ?

OpenStudy (fibonaccichick666):

Why 1?

OpenStudy (mkmkasim):

would it be u?

OpenStudy (fibonaccichick666):

Can you tell me why? or cite your source?

OpenStudy (mkmkasim):

i just remember that the derivative of arctan is 1/1+x^2

OpenStudy (fibonaccichick666):

That would be inverse tan of x, I'm asking inverse tan of x/a

OpenStudy (mkmkasim):

1/((x/a)^2 +1) ?

OpenStudy (fibonaccichick666):

Take a look up top at satelite's answer wayyyy up there

OpenStudy (mkmkasim):

so a u sub for the 1/u, and then the derivative of arctan(u/a) ?

OpenStudy (fibonaccichick666):

can you write that out for me.

OpenStudy (mkmkasim):

well l/u is the ln|u| right?

OpenStudy (fibonaccichick666):

no, you're thinking too far ahead

OpenStudy (fibonaccichick666):

just rewrite, the integral for me

OpenStudy (mkmkasim):

∫(x+4)/(u^2 +4)

OpenStudy (fibonaccichick666):

ok, so not quite. You are going to have \[\int \frac{x+4}{(x+1)^2+4} dx\] Dp you see how I got that?